OpenStudy (anonymous):

19. You are standing on a bridge over a small river, holding a stone 20 feet above the water. a. You release the stone. How long will it take the stone to hit the water? b. You take another stone and toss it straight up with an initial velocity of 30 feet per second. How long will it take the stone to hit the water? c. If you throw a stone straight up into the air with an initial velocity of 50 feet per second, could the stone reach a height of 60 feet above the water?

OpenStudy (anonymous):

One-dimensional kinematics. \(\large y=y_0+vt+0.5at^2\) y_o = initial height v = velocity a = acceleration of Earth gravity near the surface = -32.2 ft/s/s

OpenStudy (anonymous):

a. You release the stone. How long will it take the stone to hit the water? how do you solve that... there is no values..

OpenStudy (anonymous):

if we release it will undergo free fall so u=0 a=9.8 s=20 ft and find t

OpenStudy (mayankdevnani):

t=2 seconds

OpenStudy (mayankdevnani):

s=ut+1/2at^2 t=2 s

OpenStudy (mayankdevnani):

where s=20 u=0 a=9.8

OpenStudy (mayankdevnani):